show that:tan(x+a)/tan(x-a) cannot lie between the values tan square(pie/4-a) and tan square(pie/4+a)

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear ravi

let y  = tan(x+a) /tan(x-a)

        =tan{(x-∏/4) +(∏/4 +a)} / tan{( x-∏/4) + (∏/4-a)}

let tan(x-∏/4) =X

 tan(∏/4 +a) =c

tan(∏/4 -a) =d

 

now open the expression

y =[(x+c)/(1-xc)] * [(1-xd)/(x+d)]

cross multiply and simplyfy

x2(d-yc) +x(y-1)(1-cd) +(yd-c) =0

for this function have real value discrimnant must be greater than zero

  D≥0

put the discrimnant greater than zero you will get the desired result

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on

askiitians forum and become an Elite Expert League askiitian.